) Determine the equilibrium price and quantity of the above market when j=100.
Qd= Qs
given the following equations: 400 - 2P + 4j = - 100 + 3P and
400+400+100 = 3P+2P calculate price and quantity. (5)
d) Plot a graph showing the above res
1
Expert's answer
2021-08-24T14:08:13-0400

Given

"We=400-2p+4j\\Qs=100+3p,\space p=\frac{Q-100}{3}\\j=100\\Qd=400-2p+4(100)\\Qd=800-2p\\p=\frac{800-Q}{2}=400-0.5Q"

At equilibrium Qd=Qs

"100+3p=400-2p+400\\p=\frac{800-100}{5}\\=140\\Q=100+3(140)=520\space units"